### v1

# The impact of abstract and object tags on image privacy classification

Abstract
Object tags denote concrete entities and are central to many computer vision tasks, whereas abstract tags capture higher-level information, which is relevant for tasks that require a contextual, potentially subjective scene understanding. Object and abstract tags extracted from images also facilitate interpretability. In this paper, we explore which type of tags is more suitable for the context-dependent and inherently subjective task of image privacy. While object tags are generally used for privacy classification, we show that abstract tags are more effective when the tag budget is limited. Conversely, when a larger number of tags per image is available, object-related information is as useful. We believe that these findings will guide future research in developing more accurate image privacy classifiers, informed by the role of tag types and quantity.
Index Terms—  Abstractness, tags, privacy, image classification
1 Introduction
Textual information extracted from images by deep learning models (the so-called deep tags) is commonly used as a natural language description of image content to create interpretable image classifiers [1, 2, 3], and improve the performance of vision classifiers [4, 5, 6, 7, 8, 9, 10, 11]. Object tags refer to specific entities (e.g. a passport), whereas abstract tags define something that cannot be experienced directly, such as actions (e.g. running), qualities (e.g. old), and higher-level notions (e.g. spirituality, responsibility, justice) [12]. For object-guided vision tasks, the types of textual information that effectively capture visual concepts are relatively well understood [13, 7]. In contrast, limited research has examined which tag types are most suitable for addressing ambiguous and subjective tasks, such as privacy classification [14].
Most of the image privacy classifiers rely on concrete (i.e. object) tags such as ImageNet [15] class names or scene information [4, 6, 16, 9, 10]. However, since humans rely on both concrete and abstract concepts when making decisions [17] and image privacy classification is a highly subjective task [18] that requires complex human-like understanding of the visual content [19], utilising only object information for image description might be limiting for complete and deep understanding of the task. Alternatively, previous research [2] showed that the image classifier built only on eight privacy-specific concepts, which combined abstract information about spoofed, medical, violent, adult or racy content present in the image, together with information about objects (people), can achieve comparable performance to larger and more complex approaches. While the work has covered only eight tags, it suggests that privacy can be well described through a combination of concrete and abstract information.
Previous works have demonstrated that people associate diverse words with both abstract and concrete concepts when annotating images [20]. However, very little research has examined how the abstractness of tags influences the performance of classifiers, specifically for image privacy. The most relevant work to ours [21] showed that the image privacy classifier achieves the best performance when taking as input the combination of abstract and concrete tags. Yet, this work utilised tags that users assigned to images when posting them on social media. However, user tags are unstructured, sparse and are rarely available [22], which restricted the previous discussion of abstractness [21] to a single dataset, with only a small number of tags per image. Moreover, user tags significantly differ from automatically generated deep tags [4, 22], thereby limiting the scalability and applicability of these findings to future research.
In this paper, we study the impact of the tag abstractness on image privacy classification. We investigate how the type and level of detail of deep tags influence the performance of image classifiers across multiple privacy datasets. We identify a relationship between the subjectivity of the classification task and the informativeness of abstract tags for it, and provide recommendations on which concept types should be utilised in tag-based image privacy classifiers to improve the performance and understanding of privacy. We ensure a fair comparison between different tag types (concrete vs. abstract) by constraining both the dictionary size and the number of tags used for each image, and explore datasets with both subjective and object-guided annotations. As a result, we show that abstract tags play a key role in the (subjective) task of image privacy classification, particularly when only a small number of tags is used to describe an image. In contrast, with a larger number of tags, comparable performance to the one on abstract tags can be achieved using only concrete tags or a combination of the two types. For tasks with object-guided annotation, concrete tags perform better than abstract tags, however, utilising only abstract information for such tasks does not significantly lower the performance.
2 Data
We use three public datasets, each having a different approach towards privacy annotation: PrivacyAlert [4] with the most subjective annotation, VISPR [23] with object-guided annotation [24], and DIPA2 [25], which combines both subjectivity and object guidance.
PrivacyAlert contains 6800 images with binary labels (private or public). These labels were derived by aggregating subjective judgments from three to five annotators, who were instructed to assume that the images were taken by themselves and of people they know, and to decide whether each image was clearly private, private, public, or clearly public. As a result, the labels comprise the judgment of the privacy of the whole image, which is guided by a relatively small number of annotators.
VISPR contains 22112 available images annotated with respect to 67 private attributes, where the presence of any attributes in an image signals that the image is private. A subset of these attributes concerns concrete entities such as different types of documents and vehicles, which are strongly tied to the presence of specific objects in the image. The majority of VISPR attributes relate to personal descriptions, life and relationships, which require the presence of a person in the image. This skews the distribution of private images towards those depicting people [24]. Overall, VISPR annotation is considerably less subjective than PrivacyAlert and can be characterised as predominantly object-guided.
DIPA2 presents 1304 images with annotations at the object level by four annotators per image. Annotators were presented with a list of detected in images privacy-related objects and their bounding boxes, and asked to rate privacy-related properties for each object. The main properties include Privacy Threatening (PT), which indicates if the annotator thinks that the object presents privacy-threatening content, and Risk Severity (RS), which marks the perceived severity of the privacy threat posed by an object [^1] If PT=0, RS was not annotated; therefore, in such cases, we assign RS a zero value.. Since the dataset does not supply binary privacy labels at the image level, we constructed them as follows: 1) for each image, we took the highest PT and RS scores assigned by each annotator across all objects in the image, and averaged these values across all four annotators for PT and RS independently. This procedure resulted in each image receiving two aggregated annotations: PT $\in[0,1]$ and RS $\in[0,7]$. 2) In order to be on the more protective side, we assigned a final binary public label to an image only if the average PT and RS values were not higher than the midpoint thresholds of 0.5 and 3.5, respectively. All the other images were labelled as private.
As a result, the binary labels received for DIPA2 images are object-guided, as annotators judged privacy risks independently for each object, and the annotations of the most privacy-threatening objects were then propagated into image-level labels. However, unlike VISPR, where the mere presence of an attribute determines privacy, in DIPA2, the privacy annotation of each object is done independently and judged subjectively by each annotator.
3 Tag concreteness
In this section, we introduce the tags chosen for the study and describe how the degree of concreteness (which we treat as the opposite end of the same scale as abstractness) is defined for each tag.
Tag extraction. To obtain a wide range of descriptive tags for each image, we employ the commercial classifier ClarifAI [^2] https://clarifai.com/clarifai/main/models/general-image-recognition. This system has been successfully utilised in the newest interpretable, tag-based privacy classifier, achieving high performance [1]. ClarifAI generates a wide range of tags providing varied information about the image content that covers objects (e.g. ball, man, table, skin), actions (e.g. laughing, dancing), properties of the objects (e.g. pretty, dark), emotions, and abstract concepts (e.g. love, religion, wedding), etc. Formally, ClarifAI outputs up to 200 tags for each image, selected from a fixed dictionary $N=6568$-word dictionary $T=\{t_{1},...,t_{N}\}$. Each tag $t_{j}\in T$ is associated with a probability $p(i,t_{j})$ of being present in image $i$ if outputted by ClarifAI or a zero probability $p(i,t_{j})=0$ otherwise. Hence, an image $i$ can be represented as a probability vector:
$p(i,T)=(p(i,t_{1}),...,p(i,t_{N})).$ (1)
Tag concreteness. To measure the concreteness of each tag $t_{j}\in T$, we rely on the ranking provided by Brysbaert et al. [12] for a set $C$ of 40000 tags [^3] When defining $T$, we discarded 1560 tags, which were not present in $C$, from the initial 8128 ClarifAI tags.. Each tag $t\in C$ is assigned a mean concreteness rating $c(t)\in[1,5]$, aggregated across at least 22 raters, where $1$ indicates highly abstract and $5$ indicates highly concrete. Examples of tags and their mean concreteness are: spirituality (1.07), anniversary (3.15), kiss (4.48), actress (4.54), skin (4.79), people (4.82), machine gun (5.0).
To separate tags into concrete and abstract sets, we apply a threshold $4.75$. The threshold was chosen to include the maximal number of objects and the least of other types of tags in the concrete set.  [^4] While this separation relies on human labels and has some misalignments for tags with concreteness ranking close to 4.75 (e.g. tag ’football’ is considered abstract, while ’soccer’ is concrete), we observed only a few of such cases and consider them negligible due to the overall large number of tags.. Thus, we define the abstract $A=\{t_{j}\in T|c(t_{j})<4.75\}$ and concrete $B=\{t_{j}\in T|c(t_{j})\geq 4.75\}$ concept sets. For completeness, in the subsequent experiments, we also include the combined set $T=A\cup B$, which includes both abstract and concrete tags.
4 Tag representation for classifier
The abstract $A$, concrete $B$, and combined $T$ tag sets differ in dictionary size ($|A|\gg|B|$), and all sets contain tags with highly varying detection frequencies across images in privacy datasets. However, a larger dictionary size or a greater number of tags used to describe an image can provide significantly more details about the image content and consequently give an unfair advantage to a tag type. Therefore, to ensure a fair comparison across concreteness types, we apply a two-step feature selection process for each set $S\in\{A,B,T\}$, namely discriminative selection, which restricts the dictionary size, and tag sparsity control, which ensures an equal number of tags for each image.
Discriminative selection. To equalise the dictionary sizes across all tag types, we calculate the $\chi^{2}$ score between the probabilities of each tag $t_{j}\in S$ appearing in the images from a dataset and their corresponding binary privacy labels. A higher $\chi^{2}$ score indicates a stronger correlation between the tag probability and the privacy value, suggesting that the tag is important for image privacy. We keep the top $M$ tags:
$S^{\prime}=top_{M}(\chi^{2}(S))=\{t^{\prime}_{1},...,t^{\prime}_{M}\},$ (2)
with $|S^{\prime}|=M=1000$. Dictionary sizes around $1000$ are commonly used for image privacy [4, 5, 6]. The resulting feature vector $p(i,S^{\prime})$ for each image $i$ has dimension $M$. Fig. 1 shows the distribution of tag concreteness and $\chi^{2}$ scores. It is notable that both concrete and abstract tag sets contain tags with high $\chi^{2}$ scores, indicating their potential importance for image privacy classification on corresponding datasets.
Tag sparsity control. To control the number of tags $k$ describing an image $i$, for each image we retain only the top-$k$ tags with the highest probabilities $p(i,t_{j})$ among tags in $S^{\prime}$, setting all remaining tag probabilities to zero:
$p_{k}(i,t_{j})=\Big\{\begin{array}[]{ll}p(i,t_{j}),&\mbox{if }p(i,t_{j})\in top_{k}\{p(i,S^{\prime})\}\\ 0,&\mbox{otherwise}.\end{array}$
As a result for each tag type (set $S\in\{A,B,T\}$) image $i$ is described with a feature vector
$p_{k}(i,S)=(p_{k}(i,t^{\prime}_{1}),...,p_{k}(i,t^{\prime}_{M})).$ (3)
Therefore, for each image $i$ and number of tags $k$, we get three representations: through abstract $p_{k}(i,A)$, concrete $p_{k}(i,B)$, and combined $p_{k}(i,T)$ tags. Further, we investigate how the tag type and number affect the performance of classifiers trained on the corresponding representations. The selection of tag representation, described in this section, ensures that across all tag sets, the dictionary size and the number of active tags per image are matched, so that the only factor influencing performance is the abstractness of the tags.
Fig. 1: Tag concreteness and their $\chi^{2}$ score with image privacy labels in the PrivacyAlert (a) and VISPR (b) datasets. The transparency of the dots reflects the tag frequency with which the tag appears in the dataset (the more transparent, the less frequent). We visualised the tag names of a few sample points. Both abstract and concrete tag sets have tags with high $\chi^{2}$ scores, signalling that tags of different abstractness are important for image privacy.
5 Experiments
In this section, we train simple classifiers on abstract, concrete and combined tag representations while varying the number of tags per image, and analyse how these factors affect classification performance. Further, we also examine the co-occurrence patterns of abstract and concrete tags to provide insight into the observed performance trends.
For each tag type $S\in\{A,B,T\}$, the image $i$ representation is $p_{k}(i,S)$ (Sec. 4). We explore different numbers of tags with which images are described, $k\in[1,25]$. This allows us to inspect if the behaviour of the models on tags of various concreteness changes with the increase in details about an image that these tags provide. If the image can not be described with as many as $k$ tags in at least one of the tag types, the image is disregarded from the experiments on $k$ tags. For $k=25$, only 5% of images were disregarded from the PrivacyAlert and DIPA2 datasets (equally proportionally for train, val and test splits). However, for VISPR with $k=25$, 18% of images in the dataset were discarded, and for $k=18$, around 5%.
5.1 Model trained on different types of tags
Fig. 2: The achieved F1-macro (mean and std across 10 seeds) for MLP tag-based classifiers taking as input tags of different types: abstract, concrete and combined, with a varying number of tags used to describe an image. The performance is reported on three datasets: PrivacyAlert with subjective annotation, VISPR with object-guided annotation, and DIPA2 with subjective object-guided annotation. The presented F1-macro ranges vary in scale for each dataset.
In order to perform further analysis, we first train a model taking as input tag-based feature representation of images $p_{k}(i,S)$ for different types $S$ and numbers of tags $k$. We chose a simple two-layered MLP with two hidden layers [128, 32] and a ReLU activation, as the model architecture that will be used in the analysis, similar to the MLPs used in previous research [10, 2].
For each dataset, hyperparameters were selected for $p_{14}(i,T)$ inputs, representing 14 tags describing each image with a combination of concrete and abstract tags, selected on 30 runs of Optuna [^5] https://optuna.org. Further, the same hyperparameters were used to train the models on this dataset for all numbers $k$ and types $S$ of tags. The performance is recorded across 10 seeds for each model.
Fig. 2 shows the F1-macro of MLPs on concrete, abstract and combined tag image descriptions with varying levels of detail. For the subjective PrivacyAlert dataset, when the number of tags describing each image is small ($k\leq 10$), the model that takes as input abstract tags performs significantly better than the model on concrete tags by 2.80 p.p. on average across all $k\leq 10$, and with 4.09 p.p. average difference for $k\leq 5$. Additionally, the classifier taking as input 5 abstract tags performs as well as a classifier on 11 concrete tags. This behaviour on PrivacyAlert likely arises because abstract tags latently represent more contextually complex content, which helps better differentiate privacy, while the contribution of a small number of objects to privacy is ambiguous without additional context. However, when the number of tags per image exceeds $10$, the models on all tag types converge to a similar performance.
At the same time, for the object-related VISPR dataset, the performance of models on all three tag types is similar from the beginning, with a model on abstract tags on VISPR slightly underperforming by around 0.67 p.p. on average on 3 to 6 tags. Additionally, even on $k=1$ tag, the performance of the models on VISPR is only 8.37 p.p. lower than their performance on 25 tags, while in PrivacyAlert, this gap is 14.50 p.p. for concrete tags, which can be explained by the ’easier’ nature of the VISPR classification task.
On DIPA2 classifiers on abstract, concrete, and combined tag types perform more similarly than in PrivacyAlert. Additionally, on DIPA2, which combines object annotation with subjective perception of privacy of these objects, the best performance for the small $k$, especially for $k\leq 10$, is achieved by tags that combine concrete and abstract. This implies a relation between the subjectivity of labels and abstract concepts. The performance of models on DIPA2 is also characterised by a large standard deviation due to the instability of performance resulting from the limited number of images in the DIPA2 dataset.
We can see that for all three datasets, the performance increases with the number of tags, while in some cases, enlarging the description by five tags can bring up to 5 p.p. improvements, which highlights the need for detailed and lengthy image descriptions when developing classifiers and when explaining the decision of privacy. Additionally, when the number of tags is sufficiently large, $k\geq 13$ for all datasets, the performance of models on all the types of tags becomes very similar, signalling that they all convey related information for images.
5.2 Analysis of tags co-occurrence
Previously, we observed that when using a large number of tags, the model’s performance was very similar for both concrete and abstract tags. To examine whether this effect arises from the direct co-occurrence of abstract and concrete tags, we compute the Jaccard index between every pair of abstract and concrete tags within each dataset. Specifically, we use feature vectors constructed for abstract $p_{25}(i,A)$ and concrete $p_{25}(i,B)$ tag types with $k=25$ tags retained for each image. For each pair of tags $t_{a}\in A^{\prime}$ and $t_{b}\in B^{\prime}$, where $|A^{\prime}|=|B^{\prime}|=1000$, the Jaccard index is defined as
$J(t_{a},t_{b})=\frac{\sum_{[i\in D_{train}]}(\textbf{I}(i,t_{a})*\textbf{I}(i,t_{b}))}{\sum_{[i\in D_{train}]}(\textbf{I}(i,t_{a})+\textbf{I}(i,t_{b})-\textbf{I}(i,t_{a})*\textbf{I}(i,t_{b}))}$
, where $\textbf{I}(i,t)=\textbf{I}_{[p_{25}(i,t)>0]}$ indicates if $t$ is detected in image $i$ from the train subset of the dataset $D_{train}$.
Table 1: Abstract-concrete tag pairs that have Jaccard index >0.5>0.5 in image representation for each dataset.
Dataset # pairs Examples
PrivacyAlert 120 entrepreneur - cashier, automotive - mustang, smoker - ashtray, sexy - lingerie
VISPR 99 songbird - sparrow, pollination - bumblebee, adult - man, architecture - house
DIPA2 361 manicure - fingernail, craftsmanship - sewing machine, herdsman - golden retriever, glacier - ski resort
Key – # pairs: number of abstract-concrete pairs
Table 1 reports the number of abstract-concrete tag pairs that have a Jaccard index greater than $0.5$, indicating strong co-occurrence, along with the examples. For VISPR and PrivacyAlert, only around 100 concrete tags out of 1000 have a highly co-occurring abstract tag (out of 1000), suggesting that direct co-ocurrence between concrete and abstract tags is not common. For DIPA2, this number of highly co-occurring pairs is higher, 352. However, due to the small size of this dataset, 54 of these pairs involve concepts that occur only once in the entire train set. In summary, these findings indicate that while abstract and concrete tags rarely exhibit strong direct co-occurrence, when an image is described with many tags, they begin to convey similar information about image privacy (Sec 5.1).
6 Conclusion
We analysed the importance of extracting diverse content tags for image privacy classification. While previous studies suggested that abstract tags better represent images for privacy [21], we found that the tag concreteness has a substantial impact only when a small number of tags per image description are used and the annotation instructions were open (as opposed to object-guided). We demonstrated that as the number of tags increases beyond 13, the performance gap between models trained on different tag types diminishes, and they begin to perform similarly. When only a small number of tags are available, the optimal choice of tag type becomes critical and is closely related to the subjectivity of the task: for object-guided annotation, concrete tags perform best, although abstract tags remain competitive; for subjective annotation, abstract tags provide richer information and outperform concrete ones.
Considering that tag-based or interpretable classification often explains decisions with a small number of tags, we encourage future research to incorporate not only object-oriented concepts but also abstract concepts, particularly for subjective tasks, in order to capture richer and more diverse descriptions. However, since the detection of abstract tags may be more difficult [26], with the large tag budget, such tags can be successfully replaced with concrete, object-related tags without a significant impact on the performance. Furthermore, classifier performance can benefit significantly from even modest increases in the detailedness of tag descriptions, especially for tasks where labelling is not strictly object-guided.